A fully charged defibrillator contains 1.24 kJ of energy stored in a 1.06x10-4 F capacitor. In a discharge through the patient, 588 J of electrical energy are delivered in 3.12 ms. Find the voltage (i.e. potential difference) needed to store this energy in the unit.

In the above question, what average power is delivered to the patient?
Answer #1

average power delivered

P = W/t = 588 / 0.00312

P = 1.88* 10^5 W

======

Comment before rate in case any doubt, will reply for sure.. goodluck